ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: not given Gomenyuk, Sergey, Zaporozhye State University, Ukaine, Tolok, A. V., Zaporozhye State University, Ukaine, Tolok V. A., Zaporozhye State University, Ukaine The instrumental system of analysis of problems of mechanics Abstract per E-Mail in Russisch